OpenStudy (anonymous):

HELP!! sqaure root of x-3 + square root of x+2 = 5

OpenStudy (anonymous):

\[\sqrt{x-3}+\sqrt{x+2}=5\]

OpenStudy (anonymous):

\[x \approx 6\]

OpenStudy (anonymous):

I got 13

OpenStudy (anonymous):

I squared x-3 and x+2 and 5